Every year a man is paid $500 more than the previous year. If he receives $17800 over four years, what was he paid in the first year?
Answer by farohw(175)


Let x + 500 represent the first year.

Since a man is paid $500 more than the previous year, for 4 years we write:
(x + 500) + (x + 500 + 500) + (x + 500 + 500 + 500) + (x + 500 + 500 + 500 + 500) = 17,800

By adding the equation above, 4x + 5000 = 17,800
4x = 17,800 - 5000
4x = 12,800
x = 3,200

So the amount a man was paid in the first year is 3,200 + 500 = 3, 700